OpenStudy (anonymous):

solve the proportion and check your solution c+2/45=8/5

OpenStudy (anonymous):

Do you mean (c+2)/45 = 8/5?

OpenStudy (anonymous):

\[\frac{ c+2 }{ 45 }=\frac{ 8}{ 5 }\]

OpenStudy (anonymous):

like this*

OpenStudy (anonymous):

@BluFoot

OpenStudy (anonymous):

 Cross multiply. Then you get (c+2) * 5 = 45 * 8